# DREAM: Domain-aware Reasoning for Efficient Autonomous Underwater Monitoring

### v1

@@ I Introduction
Deploying Remotely Operated Vehicles (ROVs) offers a safer and more practical solution but can be challenging in uncharted underwater environments and inefficient for long-term monitoring [4]. ROV operators often need to collaborate closely with ocean scientists, but differences in background and expertise can complicate communication. Furthermore, the survival of vital ecosystems requires long-term continuous monitoring, which can make the entire operation prohibitively expensive due to the costs associated with logistics, specialized personnel, and equipment [5]. Vision-based monitoring methods have shown promise in specific domains, with applications in surveying coral reefs [6], shipwrecks [7], and oysters [8]. However, these methods are inherently imitation learning-based and have no feedback from the outer world. Recent systems such as OceanPlan [9], OceanChat [10], and AquaChat++ [11] advance natural language piloting and multi-robot coordination, but still lack robust reasoning, adaptability, and persistent spatial memory.
@@ II Related Work
Large Language Models have revolutionized the field of robotics as a source of commonsense reasoning that bridges multimodal input from the environment, such as textual instructions or visual observations, to high-level actions [13, 14, 15, 16, 17, 18, 19]. These actions are often coupled with low-level controllers [15] or robot-specific APIs [20] to yield explicit robot control commands and hence entire actionable motion plans. In our work, we leverage visual observations to generate high-level robot actions and bridge simulation to reality by generating control commands through an integrated PD controller. While our system yields high-level actions in simulation, our approach is modular and can be combined with complementary work that performs obstacle avoidance and local navigation in underwater environments [21, 22].
Underwater robot navigation has been studied extensively in the literature[23, 24] with a rich line of work focusing on tackling localization challenges [25] posed by underwater environments [26, 27, 28, 8, 29]. Many approaches have been developed for underwater monitoring and exploration, with applications ranging from cave exploration [30] to oyster detection [31]. However, they lack the idea of the feedback loop that our approach introduces to account for efficient exploration and replanning.
@@ VI Conclusion and Future Work
For future work, we plan to leverage DINOv3 [43] features as input to improve generalizability, expand the map size and environmental complexity to approximate long-term deployments, and increase the number of planning steps. Furthermore, we also intend to explore using additional models and conducting more rigorous evaluations to assess the performance of our system. Last but no least, we aim to extend our framework to 3D exploration in cave-like environments, where irregular structures and constrained visibility present additional challenges for autonomous navigation.

### v2

@@ I Introduction
Deploying Remotely Operated Vehicles (ROVs) offers a safer and more practical solution but can be challenging in uncharted underwater environments and inefficient for long-term monitoring [22]. ROV operators often need to collaborate closely with ocean scientists, but differences in background and expertise can complicate communication. Furthermore, the survival of vital ecosystems requires long-term continuous monitoring, which can make the entire operation prohibitively expensive due to the costs associated with logistics, specialized personnel, and equipment [33]. Vision-based monitoring methods have shown promise in specific domains, with applications in surveying coral reefs [28], shipwrecks [19], and oysters [25]. However, these methods are inherently imitation learning-based and have no feedback from the outer world. Recent systems such as OceanPlan [56], OceanChat [55], and AquaChat++ [39] advance natural language piloting, and Xu et al. [53, 49, 51] contribute multiple works on multi-AUV decision making, coordination, and adaptive control, but still lack robust reasoning, adaptability, and persistent spatial memory.
@@ II Related Work
Large Language Models have revolutionized the field of robotics as a source of commonsense reasoning that bridges multimodal input from the environment, such as textual instructions or visual observations, to high-level actions [13, 12, 23, 30, 11, 6, 34, 42]. These actions are often coupled with low-level controllers [23] or robot-specific APIs [31] to yield explicit robot control commands and hence entire actionable motion plans. In our work, we leverage visual observations to generate high-level robot actions and bridge simulation to reality by generating control commands through an integrated PD controller. While our system yields high-level actions in simulation, our approach is modular and can be combined with complementary work that performs obstacle avoidance and local navigation in underwater environments [29, 54].
Underwater robot navigation has been studied extensively in the literature[9, 46] with a rich line of work focusing on tackling localization challenges [14] posed by underwater environments [16, 52, 15, 25, 47, 48]. Many approaches have been developed for underwater monitoring and exploration, with applications ranging from cave exploration [10, 7] to oyster detection [26, 27, 20, 43] However, they lack the idea of the feedback loop that our approach introduces to account for efficient exploration and replanning.
@@ VI Conclusion and Future Work
For future work, we plan to leverage DINOv3 [40] features as input to improve generalizability, expand the map size and environmental complexity to approximate long-term deployments, and increase the number of planning steps. Furthermore, we intend to train a compact local VLA model based on Mamba[1] and apply pruning[17, 18] and extreme quantization[41] to enable efficient on-device inference, along with more rigorous evaluations to assess system robustness. Last but no least, we aim to extend our framework to 3D exploration in cave-like environments, where irregular structures and constrained visibility present additional challenges for autonomous navigation.